OpenStudy (goformit100):

Are the following statements true or false? Give reasons for your answers. (i) Every whole number is a natural number. (ii) Every integer is a rational number. (iii) Every rational number is an integer

OpenStudy (goformit100):

@Kitt020912 HELP ME

OpenStudy (goformit100):

@satellite73 HELP ME SIR @sami-21

OpenStudy (anonymous):

last one is false, since for example \(\frac{1}{2}\) is a rational number, but it is not an integer

OpenStudy (anonymous):

second one is true, since any integer can be written as a fraction for example \(5=\frac{5}{1}=\frac{10}{2}=...\)

OpenStudy (anonymous):

first one is a bit tricky Every whole number is a natural number it depends on where you put 0

OpenStudy (anonymous):

I don't really pay attention to term definitions, Sorry @goformit100 :l

OpenStudy (goformit100):

it's ok @Kitt020912 , and thanks @satellite73 :D

OpenStudy (anonymous):

you need to check the convention you are using for natural numbers vs whole numbers, and see if one or the other or both include 0